OpenStudy (anonymous):

1. How is the following equation a translation of f(x)=x^2 y= (x+2)^2 Translated up two units Translated down two units Translated right two units Translated left two units

OpenStudy (anonymous):

2. What is the equation of the following graph? y=2x^2-2x-1 y=2x^2-2x+1 y=2x^2+2x-1 y=2x^2+2x+1

OpenStudy (anonymous):

3. what is the equation, in standard form, of a parabola that contains the following points? (2,0) (3,2) (4,6) y=x^2+3x+2 y=x^2+3x-2 y=x^2-3x+2 y=x^2-3x-2

OpenStudy (anonymous):

One question at a time mate

OpenStudy (anonymous):

Any idea of how to do the first one?

OpenStudy (anonymous):

nope

OpenStudy (sidsiddhartha):

please post one question at a time :)

OpenStudy (anonymous):

Try putting some numbers in it and see what the output is (graph it).

OpenStudy (anonymous):

So I just graph it with random #'s then see what units it has?

OpenStudy (anonymous):

answer of que. 2 is b

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Do you know any of the others?

OpenStudy (anonymous):

ans. of que. 3 is c

OpenStudy (anonymous):

answer of que. 1 is also c

OpenStudy (sidsiddhartha):

dont just give away the answers @vipul92

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

4. A ball is thrown and it follows a parabolic path. At each horizontal distance (x) the following equation models the balls height (y) Y=-0.02x^2+0.8+3 what will the height of the ball be at a distance of 10 feet 6 7 8 9

OpenStudy (anonymous):

How did you find #2?

OpenStudy (anonymous):

first put x=0 in all equ. in which you find y=1 is right one there are two equations for it then put x=1 if you find y=1 is the right option

OpenStudy (anonymous):

que.. 4 is wrong

OpenStudy (anonymous):

Oh cool thank you. Do you know #4?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

mins it's ans. is 1.8

OpenStudy (anonymous):

There is no other way to get a different answer?

OpenStudy (anonymous):

ok then check que. once again

OpenStudy (anonymous):

\[Y=-0.02x ^{2}+0.8+3\]

OpenStudy (anonymous):

sry then i can't find it's answer

OpenStudy (anonymous):

thats fine I have 7 more if you could help with those? please @vipul92

OpenStudy (anonymous):

5. what is the standard form of the equation? y=(x+2)^2-3 y=2x^+4x-3 y=2x^+4x+7 y=2x^+4x+1 y=2x^+4x-3

OpenStudy (anonymous):

it's c

OpenStudy (anonymous):

6. Factor the following quadratic expression. X^2+14x+24 (x+8)(x+3) (x-8)(x-3) (x-12)(x-2) (x+12)(x+2)

OpenStudy (anonymous):

@vipul92

OpenStudy (anonymous):

c

OpenStudy (anonymous):

How did you solve that?

OpenStudy (anonymous):

@vipul92

OpenStudy (anonymous):

sry ans. is d first

OpenStudy (anonymous):

x^2 + 12x + 2x + 24

OpenStudy (anonymous):

x(x+12)+2(x+12) (x+12)(x+2)

OpenStudy (anonymous):

ok now do ownself

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (anonymous):

7. Factor the following quadratic expression. X^2-7x-8 (x+2)(x+4) (x-2)(x-4) (x-8)(x+1) (x+8)(x-1)

OpenStudy (anonymous):

Where do i start?